Translate the following problems into algebraic equations :
1) 42 is 40% of what number?
2) 15 is what percent of 3?
3) The sum of the page numbers on facing pages of a book is 157; what are the page numbers?

Answer by checkley77(12844)

42=.40X
X=42/.40
X=105 ANSWER.
-----------------------
15/3=5 OR 500%
------------------------
X+X+1=157
2X=157-1
2X=156
X=156/2
X=78 IS THE FIRST NUMBERED PAGE.
78+1=79 IS THE NEXT PAGE.
PROOF
78+79=157
157=157